charlotte has 3 times as many pencils in her schoolbox as Rebecca. Together they have a total of 24 pencils. How many pencils does rebecca have?
Answer by checkley71(8403)

X+3X=24
4X=24
X=24/4
X=6 PENCILS HAS REBECCA.
CHARLOTTE HAS 3*6=18 PENCILS.
PROOF:
8+16=24
24=24